OpenStudy (anonymous):

simplify the following quotient of complex numbers into the form a + bi -i/8+8i

OpenStudy (kropot72):

The form a + bi is \[0+i(8-\frac{1}{8})\] Now you can simplify it.

OpenStudy (anonymous):

multiply top and bottom by 8-8i: -i(8-8i) / (8+8i)(8-8i)=-8-8i /64 +64 = -8-8i /128 =(-1-i)/16

OpenStudy (anonymous):

In form of a + ib : = 8i - i/8 Taking i common: = i(8 - 1/8) Solving it further by taking LCM : i(64 - 1)/8 = 63i/8.. Also, it can be written as : 0 + 63i/8 Comparing it with (a + ib), a = 0 and b = 63/8.. So, in (a + ib) form : It is 0 + 63i/8 or simply 63i/8...